OpenStudy (anonymous):

0.20(60)+0.05x=0.10(60+x)

OpenStudy (anonymous):

0.20(60)+0.05x=0.10(60+x) 12 + .05x = 0.1*60+.1x 12 + .05x = 6 +.1x -6 -6 6 + .05x = 0 +.1x -0.05x -0.05x 6 = 0.05 x /.05 /.05 x = 120